(For wide-screen, it's 16/9.) Could you use to make it only one variable and find the measurements for a wide-screen TV with 40 inches as its diagonal?
Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

(For wide-screen, it's 16/9.) Could you use to make it only one variable and find the measurements for a wide-screen TV with 40 inches as its diagonal?
--------------
d^2 = W^2 + H^2
W/H = 16/9 --> W = 16H/9
d^2 = (16H/9)^2 + H^2 = 337H^2/81
d = sqrt(337)*H/9
H = 9d/sqrt(337)
------------
W = 16d/sqrt(337)
------------

Answer by stanbon(75887)

(For wide-screen, it's 16/9.) Could you use to make it only one variable and find the measurements for a wide-screen TV with 40 inches as its diagonal?
---
width = 16x
height = 9.1x
-----
Equation:
(16x)^2 + (9.1x)^2 = 40^2
------
(16^2+9.1^2)x^2 = 1600
x^2 = 4.722
x = 2.7131
---
width = 16*2.7131 = 43.41 inches
--------------
Cheers,
Stan H.
----------------